OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of 1/2.

OpenStudy (anonymous):

y − 6 =1/2 (x + 3) y = 1/2x − 6 y + 3 =1/2 (x − 6) x − 2y = 12

OpenStudy (anonymous):

check ur calculations again. that's not the answer. and what did u do in ur 2nd step?

OpenStudy (anonymous):

it has to be one of those ?

OpenStudy (anonymous):

lol! my bad! well then, just plug in the values in the standard eq....and u'll get the answer. the standard eq is y - y1 = m( x - x1 )

OpenStudy (anonymous):

y + 3 =1/2 (x − 6)

OpenStudy (anonymous):

bingo!

OpenStudy (anonymous):

@Vaidehi09 hii i am on the same problem here, can you explain to me what to do ?

OpenStudy (anonymous):

@ThatHippieKid do u know the standard eqs for line? we have used the standard eq of a line in the point-slope form here. it goes like this y - y1 = m(x - x1) here (x1,y1) is the point on the line and m is its slope. in the given problem, (x1,y1) = (6, -3) m = 1/2 just substitute these values in the standard form and u'll have ur eq! PS. really sorry for the late reply! hope i helped. if u hv any further questions, dont hesitate to ask :)